OpenStudy (anonymous):

Check if limit exists at (x,y)=(0,0). f(x,y) = (x^3y)/(x^4+y^2) .

OpenStudy (anonymous):

What have you tried? Checking numerically, the limit does seem to exist, but I'm wondering if there's a convenient analytic approach.

OpenStudy (anonymous):

Converting to polar may help: \[\begin{cases}x=r\cos t\\y=r\sin t\end{cases}~~\implies~~\lim_{(x,y)\to(0,0)}\frac{x^3y}{x^4+y^2}=\lim_{r\to0^+}\frac{r^4\cos^3t\sin t}{r^2\left(r^2\cos^4t+\sin^2t\right)}\] One can show that \(|\cos^3t\sin t|\le\dfrac{3\sqrt3}{16}\). You also have that \[|r^2\cos^4t+\sin^2t|\le r^2|\cos^4t|+|\sin^2t|\le r^2+1\] By the squeeze theorem, then, you have \[\lim_{(x,y)\to(0,0)}\frac{x^3y}{x^4+y^2}=\lim_{r\to0^+}\frac{r^4\cos^3t\sin t}{r^2\left(r^2\cos^4t+\sin^2t\right)}=\frac{3\sqrt3}{16}\lim_{r\to0^+}\frac{r^4}{r^4+r^2}\]